THE EFFECTS OF SPIN-ORBIT COUPLING ON THE SPECTROSCOPY OF JAHN-TELLER ACTIVE MOLECULES

\underline{TIMOTHY A. BARCKHOLTZ} and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, Columbus, Ohio 43210.

We have developed a program that incorporates spin-orbit coupling directly into the normal Jahn-Teller Hamiltonian. The effects of spin-orbit coupling on the electronic spectroscopy of Jahn-Teller active molecules will be presented. The inclusion of spin-orbit coupling in the Hamiltonian is a necessity when both Jahn-Teller coupling and spin-orbit coupling are significant. Furthermore, the program handles an arbitrary number of Jahn-Teller active modes. We will show that for molecules that have more than one Jahn-Teller active mode, it is necessary to perform one multi-mode Jahn-Teller calculation rather than several one-mode calculations. This combined theory has been applied to the electronic spectroscopy of the methoxy family of radicals and to the group IIA and IIB metalmonomethyl radicals. The relationship of ab initio calculations to the spin-orbit Jahn-Teller Hamiltonian will be discussed.